OpenStudy (anonymous):

find the area of a triangle with a base of 12 meters and an altitude of 14 meters?

OpenStudy (anonymous):

A = bh / 2 A = 12*14/2 now can you solve it?

OpenStudy (anonymous):

how can i do that do i multhipy

OpenStudy (anonymous):

multiply 12*14 first 12*14 is what

OpenStudy (anonymous):

168

OpenStudy (anonymous):

ok now divided 168/2 is what

OpenStudy (anonymous):

84

OpenStudy (anonymous):

there is ur answer now do you get it?

OpenStudy (anonymous):

yes thank u for your help

OpenStudy (anonymous):

ok let me ask you a questions see if you are correct

OpenStudy (anonymous):

find the area of a triangle with a base of 16 meters and an altitude of 4

OpenStudy (anonymous):

can you answer this without my help?

OpenStudy (anonymous):

whats that

OpenStudy (anonymous):

can you solve this questions to see if you understood find the area of a triangle with a base of 16 meters and an altitude of 4? Remember A + B*H/2

OpenStudy (anonymous):

sorry my bad A = b*h/2 not plus

OpenStudy (anonymous):

A=bh/2 A=16x4=64 64x4=256

OpenStudy (anonymous):

close you need 64 divided by 2 not multiply by 4 again...so the answer is 32 you doing great job.. Bye :)

OpenStudy (anonymous):

dang